### v1

# Deblurring in the Wild: A Real-World Dataset from Smartphone High-Speed Videos

Abstract
We introduce the largest real-world image deblurring dataset constructed from smartphone slow-motion videos. Using 240 frames captured over one second, we simulate realistic long-exposure blur by averaging frames to produce blurry images, while using the temporally centered frame as the sharp reference. Our dataset contains over 42,000 high-resolution blur-sharp image pairs, making it approximately 10 times larger than widely used datasets, with 8 times the amount of different scenes, including indoor and outdoor environments, with varying object and camera motions. We benchmark multiple state-of-the-art (SOTA) deblurring models on our dataset and observe significant performance degradation, highlighting the complexity and diversity of our benchmark. Our dataset serves as a challenging new benchmark to facilitate robust and generalizable deblurring models. We release the dataset and generation scripts at huggingface.
1 Introduction
Image restoration is a widely studied field in computer vision, with research focusing on restoring image quality by removing various forms of noise including blur, haze, noise, and moiréing. Image deblurring refers to the process of restoring the latent sharp image from a blurred image. Motion blur is the most common form of blur in images, and its source is diverse. Fast-moving subjects and/or camera shake can result in motion blur, as well as occlusion in motion boundaries and variation of scene depth.
Initial approaches to image deblurring [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] proposed simplified mathematical models and relied on the assumption that the blur is uniform (not varying spatially across the image) or locally linear. These approaches were formulated using the following blur model:
$B=I\otimes k+\eta$
where $I$, $B$ and $k$ denote the sharp image, the blurred image and the blur kernel respectively. $\otimes$ represents the convolution operation, and $\eta$ denotes additive noise. The restrictive blur model and assumptions of blur uniformity across image pixels make the estimation of both $I$ and $k$ from $b$ an ill-posed problem, especially when the domain consists of real-life images from various domains.
Later, machine learning based approaches have been developed for the single image deblurring problem. Earlier methods were based on convolutional neural networks (CNNs). They either involved synthesized blur images with a uniform blur kernel for training [11, 12, 13], or were trained to estimate locally linear blur kernels [14]. This is due to the absence of datasets containing real blurry images and corresponding ground truth sharp images as pairs for supervised training. More recent approaches incorporate transformers or state space models. These data-intensive approaches also often struggle to deblur images taken from real-world scenarios and suffer from a loss of generality across blurred images from all domains. For example, some methods based on generative adversarial networks (GANs) only perform well on images containing text or human faces.
The task of building a real-world dataset consisting of blurred and sharp image pairs contains multiple challenges. The primary one is ensuring the geometric and photometric alignment of the blurred image with its corresponding sharp image. The two images have to be taken at the same camera position, which is difficult because camera motion is often essential for introducing blur. There are other necessary criteria for the dataset to be realistic, such as the representation of the most common scenarios for motion blur in the dataset. The presence of noise must also be kept at a minimum level in the ground truth image.
Figure 1: Comparison of SloMoDeblur dataset with other popular datasets.
Using iPhone’s 240 fps slow-motion feature, we construct long-exposure blur by averaging frames over a duration (typically 1 second) and use the middle frame as the sharp reference. This mimics physical long-exposure photography while keeping a ground-truth reference. Our dataset comprises 50,000 blur-sharp pairs, significantly larger than existing datasets such as GoPro [15] and RealBlur [16], which contain only a few thousand samples.
2 Related Works
In this section, we discuss the current models at the frontier of image deblurring and previous image deblurring datasets developed for both inference and learning.
2.1 Image Deblurring Models
The image deblurring problem is addressed in two distinct ways. The first is blind image deblurring, where the goal is to estimate the blur kernel that caused the degradation directly. Classical methods in this domain rely on strong priors or optimization techniques to infer both latent and sharp images, under the assumption of uniform or spatially-varying blur [1, 2, 3, 4, 7, 10, 12, 9, 8, 6, 5]. The second approach involves various deep learning techniques to learn a direct mapping from a blurry image to a sharp one. Deep learning based image deblurring has seen significant advancements in recent years, with the state-of-the-art methods now achieving PSNR scores above 34 dB and SSIM values close to or exceeding 0.97. Leading approaches [17, 18] push the limits of image deblurring by leveraging adaptive reversible decoders and efficient state-space modeling, respectively. A key trend is the growing use of transformer-based architectures [19, 20, 21], which capture long-range dependencies more effectively than traditional CNNs [22, 23, 24]. Another promising direction is the integration of frequency-domain processing, as seen in models like DeepRFT+ [25] and FSNet [26], which exploit the frequency characteristics of blur. Additionally, test-time refinement techniques and lightweight models are actively researched to improve generalization and efficiency in real-world scenarios. Despite these advancements, it is challenging to handle non-uniform blur in real-world dynamic scenes and to generalize across diverse datasets. Emerging works increasingly explore self-supervised learning, Mamba-based state space models, and cross-task unification, pointing to a future where image deblurring becomes more robust and context-aware.
Several datasets have been proposed to support both learning and inference in image deblurring, and they can be broadly categorized into synthetic and real-world types. Synthetic datasets often generate blur using predefined kernels or by blending high-speed video frames, while real-world datasets aim to capture natural blur through long-exposure photography or dual-exposure setups. Many earlier datasets were primarily designed for benchmarking and kernel estimation, rather than for training deep learning models.
2.2 Synthetic Datasets
Some of the earliest synthetic deblurring datasets were designed primarily for inference and evaluation, rather than learning, due to their limited size. Levin et al. [27] introduced a dataset consisting of 32 images blurred using real camera shakes, while Sun et al. [28] proposed a synthetic dataset by convolving 80 sharp images with eight predefined blur kernels. Köhler et al. [29] further extended this idea with 48 images blurred using recorded camera motion paths to simulate spatially varying blur. While valuable for benchmarking kernel estimation methods, these datasets are too small to train deep learning models effectively. In contrast, the GoPro dataset is a large-scale synthetic dataset specifically designed for learning-based deblurring [15]. It was generated by capturing high-frame-rate video and averaging consecutive frames to simulate realistic motion blur. A similar approach was later adopted for human-centric deblurring in the HIDE dataset [30], which also uses a high-speed GoPro camera to produce synthetic blurry images of people in motion, enabling the training of models tailored to human-focused scenarios.
2.3 Real-world Datasets
Unlike synthetic datasets, real-world deblurring datasets aim to capture naturally occurring blur under authentic imaging conditions, making them more suitable for evaluating real-scene performance. Levin et al. and Kohler et al. attempted this to some extent using controlled camera motion, but these setups were still constrained and lacked scalability. Lai et al. [31] introduced a dataset of 100 real blurry images for benchmarking, but it lacked paired ground truth sharp images, limiting its utility for supervised learning. More recently, Rim et al. [16] addressed these limitations with the RealBlur dataset, which uses a hybrid camera setup to capture both long-exposure (blurred) and short-exposure (sharp) images of the same scene under natural lighting using a beam splitter. This approach produces high-quality, spatially varying real blur paired with accurate ground truth, making RealBlur one of the first real-world datasets suitable for training and evaluating deep deblurring models.
2.4 Challenges and Contributions
Despite the progress enabled by synthetic and real-world datasets, current resources remain limited in several critical aspects. Large-scale synthetic datasets like GoPro and HIDE have driven recent advances in deep learning-based deblurring, but their size and domain coverage are still constrained. GoPro contains 3,214 image pairs, while HIDE is similarly sized and focused solely on human subjects. Both are insufficient for training large modern architectures and do not generalize well to more diverse real-world scenarios.
RealBlur, one of the few real-world datasets designed for learning, improves upon earlier efforts by providing high-quality blurry-sharp pairs with better accuracy. However, it is limited in scale as well, offering fewer than 3,758 image pairs, and its data collection method introduces additional challenges. The beam splitter setup reduces the light reaching each sensor, necessitating extensive post-processing. Moreover, since the dataset was captured using DSLR cameras, the images tend to be of exceptionally high quality, failing to represent the imaging characteristics of consumer smartphones. These smartphones, with their smaller sensors and limited optics, produce more severe and complex blur due to reduced dynamic range and detail capture, conditions under which most real-world users take photos.
To address these limitations, we introduce a new large-scale dataset comprising over 42,000 blurry-sharp image pairs, more than ten times the size of the GoPro dataset, entirely using a modern iPhone. Using an averaging-based approach on high-frame-rate video, we simulate realistic motion blur in unconstrained settings, while the corresponding sharp frames serve as reliable ground truth. This scale, combined with the domain relevance of smartphone imagery, positions our dataset as a new benchmark for image deblurring. It reflects the true challenges of blur encountered in everyday photography and provides a robust foundation for developing models that generalize to the most common real-world use case: photos captured by the average user on a smartphone.
3 Dataset Construction
To construct a high-fidelity dataset for training and evaluating data-driven motion deblurring algorithms, we designed a data acquisition pipeline that captures sharp and blurred image pairs under real-world conditions at the same time. We utilize the high-speed video recording capability of the iPhone 15 Pro, a modern consumer-grade smartphone equipped with an advanced imaging sensor, wide dynamic range, and computational photography pipeline. This choice of hardware serves two critical purposes: (1) it enables dense temporal sampling of real-world motion at high spatial resolution, and (2) it ensures that the resulting dataset is well-aligned with typical consumer photography scenarios, thereby improving its generalizability and real-world relevance.
We captured video sequences at a frame rate of 240 frames per second (fps) in slow-motion mode, with each frame recorded at a resolution of 1920×1080 pixels. This configuration provides both the temporal granularity needed to model motion blur formation accurately and the spatial detail necessary for fine-grained analysis of restoration performance. Additionally, utilizing a widely available mobile device rather than a specialized camera system facilitates reproducibility and lowers the barrier for future dataset expansion or adaptation.
Figure 2: (a) Extracting pixel data from an image frame (b) Layering temporally close frames to produce a synthetic blur
3.1 Blur Formation Model
Motion blur occurs when the camera sensor integrates radiance from a dynamic scene over a non-zero exposure interval. Mathematically, the blur image $B$ is modeled as:
[…]
where $F(t)$ is the sharp image at time $t$ and $T$ is the exposure time. Since continuous integration is not feasible in digital settings, we discretize the blur as:
$B^{\prime}=\frac{1}{N}\sum_{i=0}^{N-1}F_{i},$ (2)
where $F_{i}$ is the $i$-th sharp frame sampled from high-fps video. We fixed $N=30$, resulting in an effective simulated exposure time of $T=\frac{30}{240}=\frac{1}{8}$ seconds. This simulates the blur expected in common handheld scenarios.
The image pair generation process is detailed in Algorithm 1, which creates aligned blurred and sharp images from a high-fps video sequence.
Input: Video $V$ with frames $V[0\ldots N-1]$
Output: Blurred and ground truth image pairs saved to disk
1
[…]
3 $frame\_itr\leftarrow 0$;
4
5while $frame\_itr+FRAMES\_TO\_LAYER\leq N$ do
6       $accumulator\leftarrow 0$ (empty frame);
7
8      for $i\leftarrow 0$ to $FRAMES\_TO\_LAYER-1$ do
9             $frame\leftarrow\text{float}(V[frame\_itr+i])$;
10             $accumulator\leftarrow accumulator+frame$;
11
12            if $i=\lfloor FRAMES\_TO\_LAYER/2\rfloor$ then
13                   $ground\_truth\leftarrow V[frame\_itr+i]$;
14
15
16
17      $blurred\leftarrow\text{uint8}(accumulator/FRAMES\_TO\_LAYER)$;
18
19      Save $blurred$ to images/blurred;
20
21      Save $ground\_truth$ to images/groundtruth;
22
23      $frame\_itr\leftarrow frame\_itr+FRAMES\_TO\_LAYER$;
24
Algorithm 1 Create Composite Images from Video
We iterate through the video frames (ref 1 in non-overlapping segments of 30 frames (line 1). In each iteration, a temporary variable accumulator is initialized to hold the sum of pixel values (line 1). The midpoint frame is selected as the sharp ground truth (line 1). Once all 30 frames have been accumulated, their mean is computed and converted to construct the blurred image (line 1).
3.2 Ground Truth Definition
The central frame of the accumulation window is chosen as the ground truth sharp image:
$F_{\text{GT}}=F_{\lfloor N/2\rfloor}.$ (3)
This choice minimizes temporal misalignment and is supported by the assumption that the motion within the short exposure interval is locally smooth. We assume the objects are in uniform motion with the very small temporal window, and with 240 fps, the center frame approximates a latent sharp image with negligible distortion.
3.3 Camera Sensor Quality
High-end DSLR cameras and devices like the GoPro HERO11 Black feature larger sensors and hardware-level stabilization, making them less susceptible to motion blur. Smartphones, including the iPhone 15 Pro, have smaller sensors and rely more on computational enhancements. However, this hardware limitation results in more frequent motion blur, which is representative of real-world scenarios. Thus, training models on smartphone-generated blur ensures better generalization to typical user environments.
3.4 Dataset Diversity and Real-World Relevance
One of the distinguishing features of our dataset is its diversity in both scene content and motion dynamics. We curated a broad spectrum of sequences that include various forms of motion, such as human activities (e.g., walking, running, hand gestures), object displacements (e.g., flying papers, swinging bags), and camera-induced movements (e.g., panning, shaking, tilting). The dataset captures both global and local motion patterns, including linear, rotational, and occlusive dynamics, as well as non-uniform motion blur resulting from depth variation and parallax.
Furthermore, because our dataset is collected using a widely adopted consumer device (the iPhone 15 Pro), it inherently incorporates the imaging characteristics typical of real-world photo and video capture pipelines. These include realistic noise levels, tone-mapping behavior, and in-camera processing features such as stabilization, denoising, and limited dynamic range compression. By reflecting the operational conditions under which end users typically capture photos, our dataset represents deployment scenarios better than datasets collected using professional, calibrated setups.
All video sequences and derived image pairs maintain a spatial resolution of 1920×1080 pixels, preserving consistency across samples. The dataset consists of high-quality (blurred, sharp) image pairs directly aligned in space and time. Unlike synthetic datasets that require explicit modeling of the camera response function (CRF) or the application of gamma correction, our dataset avoids such post-processing assumptions. Instead, we extract linear RGB data directly from the iPhone’s video stream, simplifying the data generation pipeline and enhancing physical realism.
The proposed dataset provides a high-quality, diverse, and practically relevant benchmark for learning-based motion deblurring. Its foundation in real-world hardware, natural motions, and precise blur synthesis makes it well-suited for training models that generalize to actual deployment settings, including mobile and consumer photography applications.
4 Experiments and Evaluation
This chapter presents a comprehensive evaluation of state-of-the-art image deblurring models on our proposed dataset. We begin by benchmarking several leading methods using standard quantitative metrics such as PSNR and SSIM on a substantial 10% test split, which includes approximately 4,200 high-resolution image pairs. The performance results are analyzed in the context of a baseline derived from direct blurry-to-ground-truth comparisons, highlighting the inherent difficulty of deblurring images from our dataset. In addition, we provide a comparative overview of existing deblurring datasets, emphasizing how our contribution stands out in terms of scale, diversity, and realism. Finally, we include qualitative visual results that illustrate model behavior across challenging real-world blur conditions, offering deeper insights into the strengths and limitations of current approaches.
4.1 Benchmarks
We report the performance of leading deblurring models on our dataset using PSNR and SSIM. This analysis focuses on how well each model handles the realistic and diverse blur patterns present in our test split. In addition to the quantitative results, we include a comparative overview of existing deblurring datasets to highlight how our benchmark differs in scale, complexity, and practical relevance.
Table 1: Performance of SOTA deblurring models on our dataset.
Model PSNR (dB) SSIM
Blurry vs. GT (baseline) 32.38 0.777
Restormer [21] 31.89 0.776
MPRNet [23] 31.98 0.785
NAFNet [32] 31.55 0.715
FFTformer [19] 32.25 0.778
DMPHN [22] 31.11 0.697
MIMO-UNet++ [24] 31.35 0.704
AdaRevD  [17] 31.46 0.75
The results shown in Table 1 show that models perform significantly worse than on traditional datasets, demonstrating the difficulty of our benchmark. The average PSNR and SSIM calculated from the input blurry image and the ground truth are 32.38 and 0.777, which sets the baseline performance. None of the papers crossed the baseline PSNR, but FFTFormer exceeded SSIM than the baseline.
Table 2: Comparison of popular image deblurring datasets
Dataset Total Pairs Train/Test Scenes Res. Type
GoPro [15] 3,214 2,103 / 1,111 33 $1280{\times}720$ Synth.
HIDE [30] 8,422 6,397 / 2,025 31 $1280{\times}720$ Synth. (human)
RealBlur [16] 3,758 3,058 / 700 110 $1280{\times}720$ Real
Levin et al. [27] 32 4 $255{\times}255$ Real (ctrl.)
Köhler et al. [29] 48 38 / 10 4 $800{\times}800$ Real (ctrl.)
Sun et al. [28] 640 80 $640{\times}480$ Synth.
Lai et al. [31] 100 blurry 100 $1600{\times}1088$ Real (unpaired)
Ours 42,045 37,841 / 4,204 843 $1920{\times}1080$ Synth. (phone)
Table 2 situates our dataset in the context of existing deblurring benchmarks. Compared to widely used datasets like GoPro and HIDE, our dataset is significantly larger in both the number of image pairs and scene diversity. With over 42,000 pairs across 843 distinct scenes, and a resolution of $1920{\times}1080$, our dataset enables high-fidelity modeling of motion blur. Notably, while it is synthetically generated via frame averaging from high-frame-rate video, the use of consumer-grade smartphone footage provides realistic and diverse blur patterns that reflect real-world camera motion. This combination of scale, diversity, and realism distinguishes our dataset from others, particularly the small-scale, controlled blur datasets like Levin et al. and Köhler et al., and the relatively lower resolution and limited scenes of older benchmarks. Our contribution fills a critical gap in the deblurring literature by offering a modern, challenging dataset that better represents deployment-time scenarios.
4.2 Visual Examples
To complement our quantitative evaluation, we present a selection of qualitative examples that demonstrate the diversity and complexity of our dataset.
Figure 3: Paired images with sharp and blurred versions.
These samples offer a visual understanding of the types of motion blur captured, as well as the realism of the scenes. By showcasing a range of scenarios, including camera shake, fast-moving objects, and human activity, we aim to illustrate the practical challenges faced by current deblurring models.
Figure  3 showcases a diverse subset of blurry-sharp image pairs from our dataset, highlighting the variety and realism of the captured scenes. The examples include a wide range of motion types such as object motion, human activity, and camera shake. It also captures a range of lighting conditions and depths of field. These samples illustrate the presence of complex and spatially varying lighting conditions often found in real-world scenarios.
5 Conclusion
In this work, we have introduced a novel large-scale real-world image deblurring dataset constructed from high-frame-rate smartphone video, addressing critical limitations of existing benchmarks in scale, realism, and domain relevance. By leveraging the iPhone 15 Pro’s 240 fps slow-motion capture, we synthesized realistic motion blur through temporal averaging and paired each blurred image with a temporally aligned sharp ground truth. This methodology yielded over 42,000 high-resolution blur-sharp pairs encompassing diverse scenes and motion patterns, significantly surpassing the scale of prior datasets.
Our comprehensive benchmarking of state-of-the-art deblurring models, including transformer-based architectures and frequency-domain approaches, revealed substantial performance degradation when evaluated on our dataset. This underscores the dataset’s complexity and reflects the challenges it poses to existing deblurring techniques, highlighting the need for more robust and generalizable models.
5.1 Limitations
While our dataset represents a significant step forward in realistic motion deblurring benchmarks, it is not without limitations. First, all scenes were recorded under well-lit conditions, primarily during daylight or in well-lit indoor environments, which means low-light scenarios are underrepresented despite being common in real-world photography. Additionally, all benchmark results were obtained using publicly available pretrained models without fine-tuning on our dataset; while this highlights generalization gaps, it may not reflect each model’s full potential performance on our data. The dataset also lacks diversity in capture devices, as all videos were recorded using a single smartphone model. This may limit the generalizability of results to other hardware with different optical or sensor characteristics. Finally, although we report standard quantitative metrics like PSNR and SSIM, we do not include a human perceptual evaluation, which would provide complementary insights into real-world deblurring quality, particularly in complex or ambiguous cases where numerical scores may not correlate with subjective preference.
5.2 Future Work
Building on the strengths of our dataset, several directions can be pursued to enhance its scope and applicability. One natural extension is to incorporate low-light and nighttime capture conditions, which would expose models to more challenging scenarios involving noise, reduced dynamic range, and sensor non-linearities. Adding data from a broader range of devices including lower-end smartphones, tablets, and action cameras could further diversify the imaging characteristics and improve model robustness across hardware variations. Another promising avenue is to augment the dataset with precise motion metadata, such as gyroscope or optical flow data, which can enable the development of motion-aware or physics-informed deblurring models. Furthermore, collecting user ratings or conducting perceptual studies would help complement traditional metrics like PSNR and SSIM with human-centered evaluations, providing a more comprehensive understanding of visual quality. Finally, leveraging our dataset for training generative or diffusion-based models could open new paths for perceptually grounded and temporally consistent deblurring performance.
By providing a dataset that closely mirrors real-world conditions encountered in everyday smartphone photography, we offer a valuable resource for developing and evaluating advanced deblurring algorithms. We anticipate that this dataset will catalyze further research into models that can effectively handle the intricacies of real-world motion blur, ultimately advancing the field of image restoration.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Table 2 contains the main contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work are discussed in section 5.1.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We assume, given the small temporal window, the objects are in uniform velocity, and the central frame is a fair assumption of the sharp image 3.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The abstract contains the URL to the dataset. We also provide the scripts to generate the sharp and blurry images in the supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset is publicly available in HuggingFace as mentioned in the abstract, and the code is made available in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [N/A]
  29. Justification: We are proposing a dataset rather than conducting experiments; therefore, training and test splits, hyperparameters, and optimizer choices are not applicable. However, we did provide the test/train split details in Chapter 4 regarding our proposed dataset.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The nature of our proposed dataset does not necessitate experiments related to statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: 1 addresses the devices used and the experimental setup.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have gone through the NeurIPS code of ethics and have conformed to them.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: 2.4 contains the impact on society and the industry of smartphone photography.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The authors reserve rights to all assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Files and their metadata are uploaded to HuggingFace.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human subjects were used for the research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects were used for the research.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLMs were used for editing and grammar checks.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.